A car weighing 1200kg is moving with a velocity of 12.0 m/s to the East. Another car, located 40 meters away, has a mass of 1800kg and is moving with a velocity of 20.0 m/s to the East as well.



a. Find the position of the center of mass of the system consisting of the

two cars.

b. Find the momentum of each of the cars.

c. Find the momentum of the entire system.


1
Expert's answer
2021-12-07T09:51:30-0500

(a) "X_C=\frac{m_1x_1+m_2x_2}{m_1+m_2}=\frac{m_1\cdot(x_0+v_1t)+m_2\cdot(v_2t)}{m_1+m_2}="


"=\frac{1200\cdot(40+12t)+1800\cdot(20t)}{1200+1800}=16+16.8t" . If "t=0" then "X_C=16\ (m)" .


(b) "p_1=m_1v_1=1200\cdot12=14400\ (kg\cdot m/s)"

"p_2=m_2v_2=1800\cdot20=36000\ (kg\cdot m/s)"


(c) "p=p_1+p_2=14400+36000=50400 \ (kg\cdot m/s)"